Cosplay Unlimited sells 720 outfits per year at an average price of $164. The terms of the sale are 1/10, net 30. The discount is taken by 66 percent of customers. What is the company's accounts receivable balance? Assume 365 days per year.
Answer #1

Total sales = 720 × $164 = $118,080

Average collection period = 0.66(10) + 0.34(30) = 16.80 days

Receivables turnover = 365/16.80 = 21.72619

Accounts receivable = $118,080 / 21.72619 = $5,434.92